OpenStudy (anonymous):

Michelle plans to cover two walls of her rectangular room with wallpaper, as shown below. A roll of wallpaper is 10 meters long and 0.5 meters wide. Assuming that there is no wastage, how many rolls of wallpaper would Michelle require to cover the two walls?

OpenStudy (anonymous):

Are you going to give us a picture to look at? ("as shown below"???)

OpenStudy (anonymous):

we got 5 m^2 for area of the roll , but as Basketweave said - we need to see/know dimensions of the two rectanlges in order to help you

OpenStudy (anonymous):

side in front 6*5 = 30 for area other wall is also 6*5 = 30 so to do both walls you have to cover 60 m sq each roll is 5 m sq 60/5 = 12 # rolls needed to cover

OpenStudy (anonymous):

thank you!

OpenStudy (anonymous):

uw